OpenStudy (anonymous):

Simplify (x-2)(x+5)/ (x^2-25) A. 6/25 B. (x-1) / (x-5) C. (x-1)/ (x+5) D. -4/25

OpenStudy (anonymous):

@ELizabethL

OpenStudy (anonymous):

do you know how to factor \[x^2-25\]?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

are you in 8th or 7th grade?

OpenStudy (anonymous):

ok then time to learn it it is always the case that \[a^2-b^2=(a+b)(a-b)\] that is how you factor the difference of two squares

OpenStudy (anonymous):

@ELizabethL 7th

OpenStudy (anonymous):

no i am in eleventh in your case \[a=x,b=5\] so \[x^2-25=(x+5)(x-5)\]

OpenStudy (anonymous):

\[\frac{(x-1)(x+5)}{(x+5)(x-5)}\] cancel the common factor

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

(x-1)/ (x+5)

OpenStudy (anonymous):

btw you had a typo in the question i think you had \(x-2\) but your answer choices had \(x-1\) so i assume that it was really \(x-1\)

OpenStudy (anonymous):

Yeah it was a type sorry and thanks

OpenStudy (anonymous):

but your answer is wrong

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\frac{(x-1)\cancel{(x+5)}}{\cancel{(x+5)}(x-5)}\]

OpenStudy (anonymous):

should be \[\frac{x-1}{x-5}\]